### v1

# Extraction of classical ergotropy

Abstract
Finding the time dependent perturbation that extracts the maximal amount of energy (a.k.a. ergotropy) from a thermally isolated quantum system is a central, solved, problem in quantum thermodynamics. Notably, the same problem has been long studied for classical systems as well, e.g., in the field of plasma physics, but a general solution is still missing there. By building on the analogy with the quantum solution, we provide the classical ergotropy extraction driving: it consists of an instantaneous quench followed by an adiabatic return. We illustrate how the solution, which is valid under an ergodic assumption, is instrumental to finding the ergotropy extracting driving in more general cases. We also show that, just like in the quantum case, the classical ergotropy splits into a coherent and an incoherent part. The presented results open new ways for practical energy recovery in the classical regime while suggesting that there is nothing genuinely quantum in the quantum ergotropy problem.
What is the maximal amount of energy that can be extracted from a thermally isolated system by means of a cyclical external driving? And what is the driving that achieves such maximal energy extraction?
These questions are in the limelight of current investigation in the field of quantum thermodynamics [1, 2, 3, 4] where they naturally emerge in the study of quantum batteries [5, 6, 7, 8], quantum heat engines [9, 10, 11, 12, 13] and dynamic cooling [14, 15, 16, 17]. While addressing these questions for quantum systems has a long history [18] they took momentum only with the work of Allahverdyan et al. [19], who coined the expression “ergotropy” for the maximal energy extractable from a quantum thermally isolated system, and provided its analytical expression along with the expression of the according "ergotropy extracting" unitary operator, see Eqs. (9,13) below.
Most notably the ergotropy extraction problem has been long studied for classical systems as well. For example, in the field of plasma physics, where, e.g., the recovery of the maximal amount of energy from a plasma of fusion products constitutes a pressing practical challenge [20, 21, 22, 23, 24].
In the pioneering and elegantly succinct work of Ref. [25] Gardner pinpointed the fundamental principles that guide the calculation of the maximal energy extraction from a classical system by means of phase space volume preserving maps, a.k.a. "Gardner free energy" [20, 21]. As pointed out recently [23] any volume preserving map can be arbitrarily well approximated by a symplectic map, therefore, formally, the Gardner free energy is indeed the maximal energy that can be extracted by a Hamiltonian flow, i.e., the classical ergotropy. Therefore, as of today, determining the classical ergotropy is a problem with a known solution.
However, the pressing practical question of what is the classical ergotropy extraction driving remains an open question. Here we first re-derive the classical ergotropy and express it in a new way, Eq. (7), and then demonstrate a general driving protocol that achieves it. The protocol is intuituve, physically motivated and is inspired by the according solution of the quantum problem whose close analogy with the classical problem is demonstrated. The presented protocol works under the provision of the ergodic hypothesis, however we shall discuss how it in fact can be employed as the basis for designing the ergotropy extraction drivings in more general cases as well. This opens new ways for the practical implementation of energy extraction protocols in the classical regime. The presented theory also sheds further light on the quantum ergotropy, showing that there is very little (if not nothing at all) genuinely quantum in it, besides the discreteness of variables.
Gardner free energy (a.k.a. classical ergotropy).– Given a classical system with unperturbed Hamiltonian $H_{0}(\mathbf{z})$ being initially in a statistical state described by the phase space distribution $\rho_{0}(\mathbf{z})$, Gardner asked what is the distribution $\rho_{1}(\mathbf{z})$ featuring the smallest energy expectation among all those that can be connected to $\rho_{0}$ by a phase-volume preserving map [25]. Gardner concluded that (i) $\rho_{1}(\mathbf{z})$ must be a decreasing function $g$ of the system unperturbed Hamiltonian $H_{0}(\mathbf{z})$, (ii) for any positive real number $\sigma$, the volume of phase space where $\rho_{1}>\sigma$ must be equal to that where $\rho_{0}>\sigma$  [^1] A simple aftermath of Gardner’s prescription is that one cannot extract energy from states of the form $\rho_{0}=g(H_{0})$ with $g$ a decreasing fuction, a fact that has been formally proved also in Ref. [27], see also [28] for a more accessible proof.. These prescriptions form the basis of the so-called "Gardner restacking algorithm" and allow to calculate the "Gardner free energy" and the function $g$ that determines the “Gardner ground state” $\rho_{1}=g(H_{0})$ [20, 21], a.k.a. the “passive state” relative to $H_{0}$ [27, 28]. Let
$R(\sigma)=\int d\mathbf{z}\theta[\rho_{0}(\mathbf{z})-\sigma]$ (1)
denote the measure of phase space space where $\rho_{0}>\sigma$ (the symbol $\theta[\cdot]$ denotes Heaviside step function). $R$ is a decreasing, hence invertible, function. Let
$\Omega_{0}(E)=\int d\mathbf{z}\theta[E-H_{0}(\mathbf{z})]$ (2)
[…]
$R(\sigma)=\int d\mathbf{z}\theta[g(H_{0}(\mathbf{z}))-\sigma]\,$, (3)
or, equivalently,
$R(\sigma)$ $=\int d\mathbf{z}\theta[g^{-1}(\sigma)-H_{0}(\mathbf{z})]=\Omega_{0}(g^{-1}(\sigma))\,.$ (4)
Applying $E_{0}$ to both sides of the above equation, we get
$E_{0}(R(\sigma))=g^{-1}(\sigma)\,.$ (5)
The Gardner ground state is therefore $\rho_{1}(\mathbf{z})=g(H_{0}(\mathbf{z}))$ with $g$ obtained by inverting Eq. (5)  [^2] Note that since $E_{0}$ is increasing, and $R$ is decreasing, $g^{-1}$ is decreasing, therefore $g$ is decreasing as expected.. The minimal energy reachable by volume preserving maps reads, accordingly:
$\breve{E}_{c}$ $=\int d\mathbf{z}H_{0}(\mathbf{z})g(H_{0}(\mathbf{z}))=\int_{0}^{\infty}de\omega_{0}(e)eg(e)$
$=$ $\int_{0}^{\infty}d\Phi E_{0}(\Phi)g(E_{0}(\Phi))=\int_{0}^{\infty}d\Phi E_{0}(\Phi)R^{-1}(\Phi)\,$, (6)
where in the first line we changed the integration variable from phase-space $\mathbf{z}$ to energy $e$, using the density of states $\omega_{0}(e)=\Omega_{0}^{\prime}(e)$, in the second line we first made the change of variable $e\rightarrow\Phi=\Omega_{0}(e)$, and then we used Eq. (5). Remarkably, by inverting the functions $R,\Omega_{0}$ in Eqs. (1,2) one swiftly gets $\breve{E}_{c}$, hence the Gardner free energy, that is, the classical ergotropy,
$\mathcal{E}_{c}=\int d\mathbf{z}H_{0}(\mathbf{z})\rho_{0}(\mathbf{z})-\int_{0}^{\infty}d\Phi E_{0}(\Phi)R^{-1}(\Phi)\,.$ (7)
Despite the interest of this and similar equivalent expressions [20, 21], they tell nothing about the pressing practical question of how to engineer a time dependent driving $V(\mathbf{z},t)$ such that the system state $\rho_{0}$ will indeed evolve onto the “Gardner ground state” $\rho_{1}$ (a.k.a the passive state of $\rho_{0}$) under the flow generated by the total Hamiltonian $H(\mathbf{z},t)=H_{0}(\mathbf{z})+V(\mathbf{z},t)$ in some time span $0,\tau]$.
The Quench-Adiabat protocol.– To get a handle on the above issue let’s look at the analogous quantum problem. Given a quantum system with Hamiltonian $\hat{H}_{0}$ being in a state described by the density operator $\hat{\rho}_{0}$, we ask what is the unitary evolution operator $\hat{U}$ that leads to the state $\hat{\rho}_{1}=\hat{U}\hat{\rho}_{0}\hat{U}^{\dagger}$ of lowest energy expectation
$\breve{E}_{q}=\min_{\hat{U}}\text{Tr}\,\hat{H}_{0}\hat{U}\hat{\rho}_{0}\hat{U}^{\dagger}\,.$ (8)
For Hamiltonians with discrete spectrum acting on a Hilbert space of finite dimension $d$ the above minimum is reached by the unitary operator [19]
$\hat{U}=\sum_{k=1}^{d}|e_{k}\rangle\langle r_{k}|\,$, (9)
[…]
Under the evolution $\hat{U}$ the system reaches the minimal energy state
$|\displaystyle\hat{\rho}_{1}=\hat{U}\hat{\rho}_{0}\hat{U}^{\dagger}=\sum_{k}r_{k}\ket{e_{k}}\bra{e_{k}}$ (12)
and the quantum ergotropy, i.e., the difference between the initial energy expectation $E_{0}=\text{Tr}\,\hat{H}_{0}\hat{\rho}_{0}$ and $\breve{E}_{q}$ amounts to
$\mathcal{E}_{q}=\sum_{j,k}r_{j}(|\langle r_{j}|e_{k}\rangle|^{2}-\delta_{jk})e_{k}\,$, (13)
where $\delta_{jk}$ denotes the Kronecker delta.
Note that the optimal unitary $\hat{U}$ is exactly realising the quantum analogue of Gardner prescriptions [25]: the “ground state” $\hat{\rho}_{1}$, Eq. (12) is a decreasing function of $\hat{H}_{0}$, and the number of eigenstates of initial and final density operators ($\hat{\rho}_{0}$ and $\hat{\rho}_{1}$), whose eigenvalue is less than some value $\sigma$ are the same, for any $\sigma>0$. This is so because the density operator eigenvalues are invariant under unitary evolution. In classical mechanics, the analogous condition is dictated by the fact that the phase space density is invariant under Liouville evolution.
What time-dependent perturbation $\hat{V}(t)$ should one enact so that the according Hamiltonian
$\hat{H}(t)=\hat{H}_{0}+\hat{V}(t)$ (14)
would induce the dynamical evolution $\hat{U}$? Notably the answer is not unique. From the mathematical point of view, perhaps the simplest solution is to chose $\hat{H}(t)=\hat{K}\doteq\hbar\frac{i}{\tau}\log U$, for $t\in[0,\tau]$, so that trivially one has $e^{-i\hat{K}\tau/\hbar}=\hat{U}$ in the time span $0,\tau]$. From the physical viewpoint, we note that $\hat{U}$ can be implemented by the following quench-adiabat (QA) protocol: (i) instantaneously quench at $t=0$ to
$\hat{H}_{1}=f(\hat{\rho}_{0})$ (15)
[…]
In classical mechanics one can do exactly the same QA protocol: (i) instantaneously quench to
$H_{1}(\mathbf{z})=f(\rho_{0}(\mathbf{z}))$ (16)
[…]
To prove that the QA protocol extracts the calssical ergotropy let
$P_{1}(\Omega)$ $=\int d\mathbf{z}\rho_{0}(\mathbf{z})\delta[\Omega-\Omega_{1}(H_{1}(\mathbf{z}))]$ (17)
denote the probability density of finding, at $t=0$, the system on the iso-$H_{1}$ hyper-surface that encloses the volume of phase space $\Omega$. The symbol $\delta$ denotes Dirac’s delta and
$\Omega_{1}(E)=\int d\mathbf{z}\theta(E-H_{1}(\mathbf{z}))$ (18)
denotes the volume of phase space enclosed by the hypersurface $H_{1}=E$. We have
$P_{1}(\Omega)$ $=\int d\mathbf{z}f^{-1}(H_{1}(\mathbf{z}))\delta[\Omega-\Omega_{1}(H_{1}(\mathbf{z}))]$
$=\int de\omega_{1}(e)f^{-1}(e)\delta[\Omega-\Omega_{1}(e)]$
$=\int d\Phi f^{-1}(E_{1}(\Phi))\delta[\Omega-\Phi]=f^{-1}(E_{1}(\Omega)),$ (19)
where in the first line we used $\rho_{0}=f^{-1}(H_{1})$, in the second line we changed the integration variable from phase-space $\mathbf{z}$ to energy $e$, using the density of states $\omega_{1}(e)=\Omega_{1}^{\prime}(e)$, and in the third line we made the change of variable $e\rightarrow\Phi=\Omega_{1}(e)$, with $E_{1}=\Omega_{1}^{-1}$ denoting its inverse.
Under the ergodic hypothesis, because of adiabatic invariance of the phase volume, the quantity $P_{1}(\Omega)$ is in fact equal to the probability density of finding the system on a hypersurface of constant $H_{0}$, enclosing a volume $\Omega$, at the end of the adiabatic driving [42, 28]. Therefore the final mean energy reads:
$\breve{E}_{c}=\int d\Omega P_{1}(\Omega)E_{0}(\Omega)\,.$ (20)
As shown in the appendix, the above expression is in fact identical to Gardner ground state energy Eq. (7), thus concluding the proof.
Quantum-classical analogy.– The classical ergotropy amounts to $\mathcal{E}_{c}=\int d\mathbf{z}\rho(\mathbf{z})H_{0}(\mathbf{z})-\int d\Omega P_{1}(\Omega)E_{0}(\Omega)$, or, in more symmetric form:
$\mathcal{E}_{c}=\int d\Omega[P_{0}(\Omega)-P_{1}(\Omega)]E_{0}(\Omega)\,$, (21)
where
$P_{0}(\Omega)=\int d\mathbf{z}\delta[\Omega-\Omega_{0}(H_{0}(\mathbf{z}))]\rho_{0}(\mathbf{z})$ (22)
[…]
$P_{0}(\Omega)$
$=\int d\Theta\int d\mathbf{z}\delta[\Omega-\Omega_{0}(H_{0}(\mathbf{z}))]\delta[\Theta-\Omega_{1}(H_{1}(\mathbf{z}))]f^{-1}(H_{1}(\mathbf{z}))$
$=\int d\Theta\int d\mathbf{z}\delta[\Omega-\Omega_{0}(H_{0}(\mathbf{z}))]\delta[\Theta-\Omega_{1}(H_{1}(\mathbf{z}))]f^{-1}(E_{1}(\Theta))$
$=\int d\Theta G[\Theta|\Omega]P_{1}(\Theta)\,$, (23)
where we first inserted the unit resolution, $1=\int d\Theta\delta(\Theta-y)$, and then we introduced the “overlap” between the microcanonical states of fixed $H_{0}(\mathbf{z})$ and $H_{1}(\mathbf{z})$, whose support respectively enclose the volumes $\Theta$ and $\Omega$  [^3] When parametrised via the enclosed volume variable $\Phi$, instead of the usual energy $E$, the microcanonical ensemble is naturally normalised and simply reads: $\rho(\mathbf{z};\Phi)=\delta(\Phi-\Omega(H(\mathbf{z})))$, where $\Omega(x)$ is the phase space volume of the region $H\leq x$ [28]. $G[\Theta|\Phi]$ can be interpreted as the conditional probability that the system is found on the hypersurface of constant $H_{0}$ that encloses the volume $\Omega$ given that it is homogeneously (i.e., microcanonically) distributed on the hypersurface of constant $H_{1}$ that encloses the volume $\Theta$, or viceversa. It is a special instance of the propagator $G_{0,\tau}[\Theta|\Omega]=\int d\mathbf{z}\delta[\Omega-\Omega_{0}(H_{0}(\mathbf{z}))]\delta[\Theta-\Omega_{1}(H_{1}(\mathbf{z}_{\tau}))]$ valid for a generic time dependent protocol taking $H_{0}$ to $H_{1}$ in a finite time $\tau$, where $\mathbf{z}_{\tau}$ denotes the evolved of initial condition $\mathbf{z}$ according to the protocol [41, 42, 28]. For a sudden quench it is $\tau=0$, so that $\mathbf{z}_{\tau}=\mathbf{z}$ and the expression boils down to Eq. (24). Furthermore just like the quantum transition probabilities $p_{ij}$ form a doubly stochastic matrix, similarly the propagator $G$ is doubly stochastic in the sense that $\int d\Theta G[\Theta|\Omega]=\int d\Omega G[\Theta|\Omega]=1$ [42, 28]
$G[\Theta|\Omega]=\int d\mathbf{z}\delta[\Omega-\Omega_{0}(H_{0}(\mathbf{z}))]\delta[\Theta-\Omega_{1}(H_{1}(\mathbf{z}))]$ (24)
Summing up, the classical ergotropy reads:
$\mathcal{E}=\int d\Omega d\Phi P_{1}(\Theta)(G[\Theta|\Omega]-\delta[\Omega-\Theta])E_{0}(\Omega)\,.$ (25)
Note the formal analogy with the quantum formula (13) whereby the discrete (adiabatically invariant) principal quantum numbers $k,j$ are replaced by the continuous (adiabatically invariant) "enclosed volumes" variables $\Theta,\Omega$, and the quantum overlap $p_{kj}=|\langle r_{j}|e_{k}\rangle|^{2}$ is replaced by the classical overlap $G[\Theta|\Omega]$  [^4] The analogy becomes even more evident if one notes that, for systems admitting a phase space description, the quantum transitions $p_{kj}$ can be written as the phase space overlaps between the Wigner functions of the eigenstates of $\hat{H}_{0}$ and $\hat{H}_{1}$ [49] and that (for ergodic systems), such Wigner functions approach the classical microcanonical distributions in the classical limit..
Quite remarkably, just like the quantum ergotropy, the classical ergotropy splits into a coherent and an incoherent part [45]. Let
$\mathcal{D}[\rho](\mathbf{z})=\int d\mathbf{z}^{\prime}\,\rho(\mathbf{z}^{\prime})\delta[\Omega_{0}(H_{0}(\mathbf{z}^{\prime}))-\Omega_{0}(H_{0}(\mathbf{z}))]$ (26)
[…]
$\beta\mathcal{E}_{c}^{c}[\rho]$ $=C[\rho]+D[\mathcal{P}[\mathcal{D}[\rho]||\rho_{\beta}]-D[\mathcal{P}[\rho]||\rho_{\beta}]\,$, (27)
where
$\mathcal{E}_{c}^{c}[\rho]\doteq\mathcal{E}_{c}[\rho]-\mathcal{E}_{c}^{i}[\rho]\quad\mathcal{E}_{c}^{i}[\rho]\doteq\mathcal{E}_{c}[\mathcal{D}[\rho]]$ (28)
define the coherent and incoherent parts of $\mathcal{E}_{c}$, $\mathcal{P}[\rho]$ denotes the passive state (i.e., the Gardner ground state) associated to $\rho$, $\rho_{\beta}=e^{-\beta H_{0}}/Z$ is a thermal state, $D[\cdot||\cdot]$ denotes the Kullback-Leibler divergence, and
$C[\rho]=D[\rho||\mathcal{D}[\rho]]$ (29)
quantifies the “coherence” of $\rho$, that is how much it is distinguisgable from its dephased companion $\mathcal{D}[\rho]$. Equations (27-29) read exactly like their quantum counterparts where $D$ denotes the quantum Kullback-Leibler divergence, and $\mathcal{D}$ is the quantum dephasing operator relative to $\hat{H}_{0}$ [45].
Example.–
Figure 1: Panels a-d): Phase space sketch of extraction of ergotropy from a non-stationary Gaussian state, Eq. (31) of a harmonic oscillator, Eq. (30). a) The phase density rotates under the unperturbed dynamics. b) At $t=0$ the harmonic potential gets instantaneously displaced so as to lock the state. c,d) In the time span $0,\tau\gg 1]$ the potential is slowly returned to its initial position. Panels e-f): Sketches of the states $\rho_{0}$, $\rho^{\prime}$, Eqs. (34,35). In all panels darker blue denotes higher density. The solid lines denote iso-$H_{0}$ curves in phase space
Consider a 1D Harmonic oscillator
$H_{0}(q,p)=\frac{p^{2}}{2}+\frac{q^{2}}{2}$ (30)
being at $t=0$ in the Gaussian state
$\rho_{0}(q,p)=\frac{1}{2\pi\sigma}\exp\left[-\frac{(q-q_{0})^{2}+p^{2}}{2\sigma}\right]\,.$ (31)
See the sketch in Fig. 1a). The state is clearly non-stationary (because $\{H_{0},\rho_{0}\}\neq 0$) and would evolve by rotating around the phase space origin at angular frequency $\omega=1$, without changing shape, if unperturbed. By suddenly displacing the harmonic potential minimum at $q=q_{0}$ blocks the state at the bottom of the displaced harmonic potential: the state is now stationary (in fact passive) relative to the new Hamiltonian
$H_{1}(q,p)=\frac{p^{2}}{2}+\frac{(q-q_{0})^{2}}{2}=f(\rho_{0}(\mathbf{z}))\,$, (32)
where $f(x)=-\sigma\ln(2\pi\sigma x)$ is a decreasing function, see Fig. 1b). By adiabatically moving the harmonic potential minimum back to the origin of the $q$-axis, the state evolves onto
${\rho_{1}}(q,p)=\frac{1}{2\pi\sigma}\exp\left[-\frac{(q^{2}+p^{2})}{2\sigma}\right]=\frac{e^{-H_{0}(q,p)/\sigma}}{2\pi\sigma}\,$, (33)
because each instantaneous Hamiltonian visited by the protocol is ergodic, see Fig. 1c,d). Since the state is passive relative to $H_{0}$ no more energy can be extracted therefrom.
Discussion.– There are a number of issues that limit the applicability of the QA protocol for ergotropy extraction. Let’s consider, for example, an initial state of the harmonic oscillator featuring a central valley, e.g.:
$\rho_{0}(q,p)$ $=\exp{\left[-\frac{(p^{2}+(q-q_{0})^{2})^{2}}{4}+\frac{p^{2}+(q-q_{0})^{2}}{2}\right]}/N\,$, (34)
where $N$ is the normalisation factor, see the sketch in Fig. 1e). Any choice of decrasing function $f$ would result in a Mexican hat shaped Hamiltonian $H_{1}$ which could be very hard to be implemented. Furthermore since any such $H_{1}$ would feature a central peak, not all regions $H_{1}<E$ would be simply connected, unlike those for which $H_{0}<E$, therefore they cannot be smoothly and adiabatically linked to each other. How to get around such problems? In this case one could still apply the QA protocol, but with the physical $H_{1}$ in Eq. (32), which has the same level hypersurfaces as $\rho_{0}$. Having $H_{1}$ the same topology of phase volumes as $H_{0}$, by adiabatically return to $H_{0}$, the state would be mapped onto
$\rho^{\prime}(q,p)$ $=\exp{\left[-\frac{(p^{2}+q^{2})^{2}}{4}+\frac{p^{2}+q^{2}}{2}\right]}/N=\frac{e^{-H^{2}_{0}+H_{0}}}{N}\,$, (35)
which is less energetic than $\rho_{0}$, but not yet passive although it is stationary (because it is a function of $H_{0}$, but not a decreasing one). In fact $\rho^{\prime}$ has a central valley, see the sketch in Fig. 1f). Using the quantum jargoon, one would say that the state $\rho^{\prime}$ features population inversions. In order to decrease the energy further one should permute rings of higher populations with rings of same volume with lower population and higher (average) energy. The smaller the rings thickness, the closer one can get to the ergotropy extraction, by increasing the number of such permutations. Interestingly, physical drivings that indeed achieve such permutations have been described in Refs. [47]. Notably, such permutation protocols defy the conditions of validity of the adiabatic theorem (because they aim at adiabatically connecting topologically inequivalent regions), but, as numerically demonstrated in Ref. [48], they can nonetheless approach the wanted permutation with very high accuracy. Thus the breakdown of the ergodic hypothesis does not seem to be such a dramatic problem for ergotropy extraction (as much as it is not dramatic for the mechanical foundations of thermodynamics [35, 28]).
In general it is expected that the QA protocol works well whenever $H_{0}$ and $\rho_{0}$ have topologically equivalent phase space structure. In particular no problems occur when the regions $H_{0}\leq E$ and $\rho_{0}>\sigma$ are all simply connected as is the case in the first example above.
Conclusions.– By bridging between quantum thermodynamics and classical plasma physics, the present work offers new viewpoints to both fields. On one hand it provides a guiding principle for energy extraction from classical systems based on the QA protocol (or variants thereof, e.g., as discussed above); on the other it suggests that the quantum theory of ergotropy has no genuinely quantum feature. The two pictures, classical and quantum, are fully one the close analogue of the other. Specifically a decomposition into coherent and incoherent part exists just like in the quantum case, corroborating the idea that energy coherences do not constitute a genuine quantum thermodynamic advantage [46].
Acknowledgments.– The author thanks Frank Ernesto Quintela Rodriguez for pointing out the plasma physics literature and for suggesting the relevance of the present results. Discussions with Vasco Cavina and Marcus Bonança are gratefully acknowledged. The material presented here is based on lectures delivered at the University of Pisa, Italy, in the academic year 2024/25.
I Appendix
I.1 Equivalence of Eq. (6) and (20)
Equation (19) holds regardless of the specific choice of decreasing function $f$. Choosing $f=R$, with $R$ in Eq. (1) we get $P_{1}(\Omega)=R^{-1}(E_{1}(\Omega)$). Furthermore with the specific choice $f=R$, the function $\Omega_{1}$ boils down to the identity. In fact we have
$\Omega_{1}(x)$ $=\int d\mathbf{z}\theta(x-R(\rho_{0}(\mathbf{z})))$
$=\int d\mathbf{z}\theta(\rho_{0}(\mathbf{z})-R^{-1}(x))=R(R^{-1}(x))=x\,.$ (36)
Therefore the inverse of $\Omega_{1}$ is also the identity, i.e., $E_{1}(\Omega)=\Omega$, in this specific case. Summing up
$P_{1}(\Omega)=R^{-1}(\Omega)\,$, (37)
showing that Eq. (20) coincides with Eq. (6).
I.2 Proof of Eqs. (27)
We first note that, for any function $f$, it is
$\int d\mathbf{z}\,\rho(\mathbf{z})f(H_{0}(\mathbf{z}))=\int d\mathbf{z}\mathcal{D}[\rho](\mathbf{z}))f(H_{0}(\mathbf{z}))\,.$ (38)
In fact we have:
$\int d\mathbf{z}\mathcal{D}[\rho](\mathbf{z})f(H_{0}(\mathbf{z}))$
$=\int d\mathbf{z}\int d\mathbf{z}^{\prime}\,\rho(\mathbf{z})\delta[\Omega_{0}(H_{0}(\mathbf{z}^{\prime}))-\Omega_{0}(H_{0}(\mathbf{z}))]f(H_{0}(\mathbf{z}))$
$=\int d\mathbf{z}\int de\,\omega_{0}(e)\rho(\mathbf{z})\delta[\Omega_{0}(e)-\Omega_{0}(H_{0}(\mathbf{z}))]f(H_{0}(\mathbf{z}))$
$=\int d\mathbf{z}\int d\Phi\,\rho(\mathbf{z})\delta[\Phi-\Omega_{0}(H_{0}(\mathbf{z}))]f(H_{0}(\mathbf{z}))$
$=\int d\mathbf{z}\,\rho(\mathbf{z})f(H_{0}(\mathbf{z}))\,$,
where in the second line we have used the definition (26), we then did the usual changes of variables $\mathbf{z}^{\prime}\to e\to\Phi$ (see, e.g., Eq. 6) and finally used $\int d\Phi\delta(\Phi-y)=1$.
Note that a special case of Eq. (38) is
$\int d\mathbf{z}\rho(\mathbf{z})H_{0}(\mathbf{z})=\int d\mathbf{z}\mathcal{D}[\rho(\mathbf{z}))H_{0}(\mathbf{z})\,$, (39)
saying that the dephasing operator $\mathcal{D}$ does not alter the energy expectation.
As a consequence of Eq. (38) we have:
$C[\rho]=\mathcal{H}[\mathcal{D}[\rho]]-\mathcal{H}[\rho]\,$, (40)
where $\mathcal{H}$ denotes the classical information
$\mathcal{H}[\sigma]=-\int d\mathbf{z}\,\sigma(\mathbf{z})\ln\sigma(\mathbf{z})\,.$ (41)
To see that note, from Eq. (26), that $\mathcal{D}[\rho]$ depends on $\mathbf{z}$ through $H_{0}$, Eq. (26), namely it is of the form $f(H_{0})$. Therefore
$C[\rho]$ $=D[\rho||\mathcal{D}[\rho]]$
$=\int d\mathbf{z}\rho\ln\rho-\int d\mathbf{z}\,\rho\ln\mathcal{D}[\rho]$
$=\int d\mathbf{z}\rho\ln\rho-\int d\mathbf{z}\,\mathcal{D}[\rho]\ln\mathcal{D}[\rho]\,.$ (42)
Another useful formula is:
$D[\sigma||\rho_{\beta}]=\beta\int d\mathbf{z}(\sigma(\mathbf{z})-\rho_{\beta}(\mathbf{z}))H_{0}(\mathbf{z})-\mathcal{H}[\rho]+\mathcal{H}[\rho_{\beta}]\,$, (43)
where $\rho_{\beta}$ is a thermal state:
$\rho_{\beta}(\mathbf{z})=\frac{e^{-\beta H_{0}(\mathbf{z})}}{Z_{0}},\quad\beta>0\,.$ (44)
The proof goes as follows
$D[\sigma||\rho_{\beta}]$ $=\int d\mathbf{z}\,\sigma\ln\sigma-\int d\mathbf{z}\,\sigma\ln\rho_{\beta}$
$=-\mathcal{H}[\sigma]-\int d\mathbf{z}\,\sigma\ln(-\beta H_{0}-\ln Z_{0})$
$=-\mathcal{H}[\sigma]+\beta\int d\mathbf{z}\,\sigma H_{0}-\beta\int d\mathbf{z}\,\rho_{\beta}H_{0}+\mathcal{H}[\rho_{\beta}]\,$,
where, in the last line we used the well known formula:
$\mathcal{H}[\rho_{\beta}]=\beta\int d\mathbf{z}\,\rho_{\beta}(\mathbf{z})H_{0}(\mathbf{z})+\ln Z_{0}\,.$ (45)
We are now ready to prove Eq. (27). We have:
$\beta\mathcal{E}_{c}^{c}$ $=\beta(\mathcal{E}_{c}-\mathcal{E}_{c}^{i})=\beta\int d\mathbf{z}H_{0}(\mathcal{P}[\mathcal{D}[\rho]]-\mathcal{P}[\rho])$
$=\beta\int d\mathbf{z}H_{0}(\mathcal{P}[\mathcal{D}[\rho]]-\rho_{\beta})-\beta\int d\mathbf{z}H_{0}(\mathcal{P}[\rho]-\rho_{\beta})$
$=D[\mathcal{P}[\mathcal{D}[\rho]]||\rho_{\beta}]+\mathcal{H}[\mathcal{P}[\mathcal{D}[\rho]]]-\mathcal{H}[\rho_{\beta}]$
$\phantom{xxx}-D[\mathcal{P}[\rho]||\rho_{\beta}]-\mathcal{H}[\mathcal{P}[\rho]]+\mathcal{H}[\rho_{\beta}]\,.$
Equation (27) follows by noticing that $\mathcal{H}[\mathcal{P}[\rho]]=\mathcal{H}[\mathcal{\rho}]$, $\mathcal{H}[\mathcal{P}[\mathcal{D}[\rho]]]=\mathcal{H}[\mathcal{D}[\rho]]$, because, by construction, any state (be it $\rho$ or $\mathcal{D}[\rho]$), is linked to its passive companion ($\mathcal{P}[\rho]$ or $\mathcal{P}[\mathcal{D}[\rho]]$, respectively) by a volume preserving map, and it is a well known fact that the classical information is invariant under such transformations.